OpenStudy (vortish):

The perimeter of a rectangle is the sum of its four sides and if the width is 8inches and its perimeter is 42 inches what is the length

OpenStudy (anonymous):

let the length be 'l' inches then l+l+8+8 =42 or 2l +16 =42 2l =42-16 or 2l =26 or l=26/2 or l=13 hemce length is 13 inches